A boat takes 2 hours to travel 8 km and back in still water .If the velocity of water is 4m/hr ,the time taken for going upstream 8km and coming back is.?Thanks in advance :)

Grade:11

5 Answers

Ritha
6 years ago
From the first condition we can determine the speed of boat....v = 8km / 2hr = 4km/h
...the velocity of water is 0.004km/h
During upstream motion the net velocity would be 4-0.004 = 3.996km/h
So time taken (t) = distance/speed=8 / 3.996 = 2.002hr
During downstream motion net velocity will be 4 + 0.004 = 4.004km/h
=> t = 8 / 4.004 = 1.998hr
Aanshi
6 years ago
Time=distance/velocityVbr= 8km/hVrg=4km/hLet t1 and t2 be the time taken by boat to travel in upstream and downstream respectivelyUpstream=8-4=4Downstream=8+4=12T1=8/4=2hrT2=8/12=2/3Total time=2+2/3=2 hr and 40 mins
Surendra
6 years ago
time = distance/velocity  
when in the direction of current velocity = 8+4= 12 
t1= 8/12 = 2/3 hr 
when against the current velocity 8-4 = 4 
t2= 8/4 = 2 hr. 
total time = 2+ 2/3 = 8/3 hrs. = 2 hrs&40 minutes
Vanshika
5 years ago
Here the speed of water is still. So, they will have equal velocity and equal time taken.
             ☆ Total time = 2 hrs 
                   Time for upstream = 1 hr
                                   downstream = 1 hr 
              ☆ their velocity is 8 kmph each.
So, new downstream velocity will be (8+4) = 12 kmph
       new upstream velocity will be (8-4) = 4 kmph
 
 
              TIME TAKEN = DISTANCE / VELOCITY 
                                                            (distance is 8 km in both cases )
                                      =( 8/12 + 8/4 ) hours 
                                      =( 2/3 + 2 ) hrs 
                                      = 160 mins = 2 hr 40 mins 
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
Time = distance/velocity  
when in the direction of current, relative velocity with respect to ground = 8+4= 12 
t1= 8/12 = 2/3 hr 
when against the current, velocity w.r.t ground = 8-4 = 4 
t2= 8/4 = 2 hr. 
total time = 2+ 2/3 = 8/3 hrs. = 2 hrs&40 minutes
 
Hope it helps.
Thanks and regards,
Kushagra